OpenStudy (anonymous):

Find the maximum and minimum values of the following expression for 0<=theta<360. \(\frac{ 1 }{ 6+\cos ^{2} \Theta }\) @Callisto

OpenStudy (callisto):

The smaller the denominator, the larger the value of the function, agree?

OpenStudy (anonymous):

i don't understand actually

OpenStudy (callisto):

1/2 and 1/4, which one is larger?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

ok, got it

OpenStudy (callisto):

So, to find the max, we need to minimize the denominator. Consider the range of cosθ, -1 <= cosθ <= 1. So, what is the range of (cosθ)^2?

OpenStudy (anonymous):

1<=cosθ<=1 ???!!

OpenStudy (anonymous):

let \[\frac{1}{6+\cos^2\theta}=f(\theta)\] \[f'(\theta)=\frac{\sin2\theta}{(6+\cos^2\theta )^2}=0\implies \sin 2\theta =0,2\theta=90+360k\\ \therefore \theta=45+180k,\theta=45,225\\f(45)=2/13=f(225)\]

OpenStudy (callisto):

1<=cosθ<=1 ^That means cosθ = 1... It should be 0<= cosθ <=1. Understand why?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

not sure about min

OpenStudy (anonymous):

@Callisto understand

OpenStudy (callisto):

So, you can find the min and max now. Substitute the min of (cosθ)^2 to find the max of the function; and substitute the max of (cosθ)^2 to find the min of the function.

OpenStudy (anonymous):

1/7 <= ... <= 1/6

OpenStudy (callisto):

So, do you get the answer now?

OpenStudy (anonymous):

max : 1/6 min : 1/7

OpenStudy (callisto):

Yup :)

OpenStudy (anonymous):

thanks ;D i will remember the rule